PIs for A2056: E. J. Korpela and B.-C. Koo. Project is commensal with
ZOA project A2051 - see their observing log.
Telescope took ~7 minutes to slew to starting position (had to go from
az=270 to az=100 - telescope position, not source position), so started with
Lambda 01 of scan 4 (ie. day_04). So 18 Lambdas done on this day.
ALFA Rotation angle was set to 6 degrees at start of observation.
Maximum declination reached was approx. 21:02, so should get the desired
tiny overlap with second 7-day run starting 15 Oct.
Observing went smoothly. Commensal observers today were P. Henning and
Beam tracks did not show evenly spaced beams: see Day Zero
beam positions. Need to try different angle.
Again telescope took some time to arrive at starting SFS position, and as a
result also needed some time to slew to starting basketweave position.
Began on Lambda 02 of scan 2 (day_02) - 17 Lambdas observed today.
Trying ALFA rotation angle of 8.5 degrees. Upon inspection of mh files,
it appears the beam spacing is correct BEFORE basketweave starts, but
once loops start spacing is incorrect. See?. Need to sort this out.
Observing went smoothly. Commensal observers were PH and JB again.
Began early, so managed to get Lambda 00 of scan 0 (day_00) - ie. the
very first Lambda of the program. So 19 Lambdas observed.
Using rotation angle of 0 degrees, knowing it will improve beam spacings
compared to previous two days, but still not optimal angle. Offline
investigations in preparation of x108 parallactic angle correction tests
are underway which will help pin down the optimal angle.
At start of run, discovered galfa was still taking data from TOGS the
previous night - about 4 hours' worth of useless spectra measured. I
notified Mary Putman of this issue.
Observing went well. PH and JB present for commensal observing.
Had to run start_gpib on wappserv this morning!
Had to slew from az=270, so again had to wait about 7 minutes at start
of run. Began on Lambda 02 of scan 5 (day_05). I did not start taking
data with galfa until about 5 seconds after SFS had started. So SFS
spectra may be missing one frequency. 17 Lambdas done today.
Rotation angle for ALFA is -3 degrees, and was left at this value for
the remainder of the "low dec" portion of the run (14 Oct inclusive). This
angle gave quite good beam spacings.
Commensal observer was JB.
Ran x108 tests, which are documented here.
Got an early start, so began on Lambda 00 of scan 1 (day_01). 19 Lambdas
observed today. ALFA rotation angle left at -3 degrees.
From ~5700 seconds to ~6400 seconds, galfa had a slight increase of overflow
messages (about 0615 to 0628 AST), but these were low bit error flags.
Smooth run. JB commensal. Ran another x108 test afterward.
Another early start. Immediately upon beginning taking data with galfa,
got high bit overflow messages of type A3 MLFS=3000. This message was
repeated for all 7 beams, every second of the observations.
Stopped galfa between SFS and basketweave observations, to see if I made
mistake setting levels or anything else. So first FITS file will be
small and only contains SFS portion of observation.
Restarted data taking on galfa, overflow messages continued. This was
scan 3 (day_03), and began on Lambda 00, giving 19 Lambdas observed.
Commensal observer, JB, did not see much evidence of RFI in wapps spectra.
By contrast, VNC viewer of galfa spectra showed evidence of RFI in its
wide band, mainly from about 1435 to 1450 MHz. Beam 3 looked to be
most severely affected. Narrow band doesn't look too bad - salvageable?
Ran another x108 test after A2056. RFI persistent there too.
Epilogue: P. Perillat et al. looked into problem, and about 11 hours after
end of A2056 it was discovered that another receiver's shutter was not
properly closed, and so its motor was running and causing the RFI.
Final day on lower dec portion. Meaning scan 6 (day_06), and got another
early start which allowed us to start on Lambda 00 and get all 19 Lambdas.
No RFI/overflow problems at all. Commensal observer was B. Catinella.
Whole run went very smoothly. At start of my x108 test, I accidentally
restarted A2056, so a tiny bit of SFS was redone. No big deal.
Summarizing scan order of A2056_l: 4,2,0,5,1,3,6
Update! Oct 27: Check out the lovely Beam Spacings!
First run of upper dec portion. Only 30 minutes allotted to uncover
ALFA, but Willy got it done. He called from the Gregorian to let me know
I could start, but I waited about ten minutes anyway. Began SFS at
about 0436 AST. Willy got to control room less than 2 minutes after I
started pt2 file (which starts steering the telescope).
Began on Lambda 00 of scan 5 (day_05), get 18 Lambdas in upper dec portion
(not 19 as in lower - main reason is slightly longer scan time, 187 seconds
versus 185 seconds).
Lowest dec is 20:59:57, so we get very tiny overlap with lower dec, as
desired. Commensal observer is E. Momjian.
Desired ALFA angle is -13 degrees, but noticed that angle was displayed
as 71.44 degrees on dataview. So made several unsuccessful attempts to
change angle to -13 degrees within CIMA. Tried to rotate at top or bottom
of scans, so as not to disturb data tracks. With help of EM, managed to
force angle change using ALFA motor control gui on dataview. This did
not occur until partway through downward scan on Lambda 07 (the 8th
loop, when telescope was pointing at RA 5.67, DEC 22.5). Beam tracks
are greatly improved once this change was successfully made. See if
YOU can tell where the rotation occurred.
Got a small early headstart - gave us time to move telescope into position.
Upon setting levels on galfa, noticed dac=0 for all beams and rms levels
were very low too. Willie then realised he hadn't turned the receivers
on yet, so that got done just as we were ready to start observing.
Began on Lambda 00 of scan 0 (day_00), so total of 18 Lambdas today. The
ALFA rotation angle went to -13 degrees with no troubles today.
Midway through run, began to notice small RFI peaks in galfa wideband,
appearing regularly with a period of 12 seconds. Many radars below 1400
MHz have that period, so we must have been picking those up.
Commensal observer was C. Springob, and will be for the remainder of the
run. Today's observations went very smoothly overall.
No early start today; expt. with 327 Rx trying to catch up on lost time.
As a result we began on Lambda 01 of scan 6 (day_06), giving us 17 Lambdas
for today's run. ALFA rotated to -13 degrees normally.
12-second periodic RFI spikes were again present, so CS and I submitted a
report using the CIMA gui for RFI.
From 2620 to 2650 seconds into run (during Lambda 06), noted a series of
overflow messages in beams 3 and 4 (first in 4 followed by 3). CS also
noted this happening, visualised through wideband "streaks" across the
WAPP waterfall plots. A "Whoa" moment. No other noteable occurrences
Slightly early start, which was good because telescope was stowed at 270.
Still, missed first loop and began on Lamda 01 of scan 1 (day_01). A total
of 17 Lambdas again today.
Saw same transient signal, again around 515 AST. This time it was in beams
0,1,5 and 6. It lasted about a minute - 2870 to 2930 seconds since galfa was
started. Coords were roughly RA=5h36m, DEC=22d14'. Maybe we'll see it again?
Smooth run. Norberto let me do another x108 test
No early start, as 327 observers using extra time between their slot and ours.
Still, telescope was in a good position and we began on time, starting with
Lambda 00 of scan 3 (day_03). Thus we scored 18 Lambdas today.
Terrifically smooth run. I'm starting to think the transient signal we saw
the last 2 days was the Crab Nebula (RA=5h35m, DEC=22d). It appeared the last
two mornings at the same LST, not the same AST. We didn't see any strong
evidence of it today, but scan 3 is probably basketweaving in an opposite
sense to, say, scan 6.
We were given an early start, which allowed us to start in time to easily get the
first Lambda (Lambda 00) of scan 2 (day_02). Tomorrow will be scan 4, the last
series of basketweaves for this observing run. 18 Lambdas observerd today.
Things ran smoothly as usual. We (CS and I) gathered further compelling evidence
that the transient signal we have been seeing is the Crab Nebula. It showed up
again at 635 LST, causing overflows in beams 2, 3 and 4. Beams 0 and 5 also
showed some signal in the dataview displays. This occurred from 3335 to 3375
seconds since the start of datataking on galfa. CS took some screenshots of his
Crab Nebula in Christmas colours
Crab Nebula in Heat
Crab Nebula in RGB
Crab Nebula in Ocean view
Crab Nebula in Redless and Greenless display modes.
Willy let us start early, and almost got telescope into position in time to start on
Lambda 0, but just missed it. So we got 17 Lambdas, starting on Lambda 01, for the
final run, which is scan 4 (day_04). We had to start_gpib the wapps again today.
Another very smooth run. The Crab Nebula showed up in (mostly) beam 1 again, and
S. Gibson showed up to see how we do our observations.
Data reduction is already underway! We are trying out the newest version of the
GSR (GALFA Spectral-line Reduction) software on it. The mh and lsfs files have
already been made for each day! Amazing how fast things run on the computers here
at Arecibo...this place is great.
Summarizing scan order of A2056_u: 5,0,6,1,3,2,4
Check out the high quality of the data in this figure. This
rather colourful figure
is from one of Carl Heiles's diagnostics, and the lower rows in each figure indicate
data for which the beams are operating properly. The only place where we had a little
trouble was on October 13, which if you have been reading this whole log, you'll know
that's the day we had RFI problems (the 6th day of observing). Isn't that neat?
Here's an mpeg movie of an HI cube